OpenStudy (anonymous):

Find the center of the circle with equation x^2+y^2-4x+6y+1=0 A) (-2,3) B) (2,3) C) (2,-3) D) (-2,-3)

OpenStudy (jlastino):

hey this is similar to the one yesterday xD just do the completing the square method

OpenStudy (anonymous):

i just did it the way you tought me :D i got -2,3 is that right?

OpenStudy (rogue):

Its (2, -3)

OpenStudy (jlastino):

you got the opposite signs :))

OpenStudy (jlastino):

i forgot to tell that you should equate the ones inside the parentheses to 0

OpenStudy (anonymous):

Also for this one Find the center of the ellipse with the equation 3x^2+4y^2+18x-32y=5 i got 3,-4

OpenStudy (jlastino):

it's -3 ,4

OpenStudy (jlastino):

can you show me the equation?

OpenStudy (anonymous):

i keep getting switched up

OpenStudy (jlastino):

for example you got (x-2)^2 + (y+3)^2 =4 x-2 = 0 x=2 y+3=0 y=-3

OpenStudy (anonymous):

3x^2+4y^2+18x-32y=5 (3x^2+18x)+(4y^2-32y) (6x^2+8y^2) 3,-4

OpenStudy (anonymous):

those are the steps i took ^^

OpenStudy (jlastino):

oh don't group together the x and y they should be inside different parentheses (3x^2+18x )+(4y^2-32y )=

OpenStudy (anonymous):

oh ok

OpenStudy (jlastino):

ok here's something that could also help you notice that 3 and 18 can be are multiples of 3 you can take out the three to simplify the expressions 3(x^2+6x ) + 4(y^2-8y+ ) =

OpenStudy (anonymous):

ok

OpenStudy (jlastino):

then do the same as before 3(x^2+6x+9) + 4(y^2-8y+16) = 5 +27 + 64 ( i got 27 and 64 because i multiplied them to the number outside the parentheses respectively 3 and 4)

OpenStudy (jlastino):

you can check out this website if you like to know more about equations on circles http://www.analyzemath.com/CircleEq/Tutorials.html

OpenStudy (anonymous):

ok i got it :)